OpenStudy (vladthepro):

I'm looking for... ...help :D I think it's A. Am I correct? What is the solution to the equation 9(x + 1) = 27? x = 2.5 <--- x = 0.5 x = −0.5 x = −3.5

OpenStudy (photon336):

okay so step #1 we can divide both sides by 9

OpenStudy (photon336):

hey @VladthePro this is the problem right 9(x+1) = 27

OpenStudy (vladthepro):

Sorry abt that it's 9^(x+1) = 27

OpenStudy (photon336):

oh no wonder lol

OpenStudy (photon336):

okay so far \[9 = 3^{2}\] \[27 = 3^{3}\] agree?

OpenStudy (vladthepro):

yes

OpenStudy (photon336):

If you haven't learned logs yet, you can change the bases \[3^{2*(x+1)} = 3^{3}\]

OpenStudy (photon336):

so that the bases are both 3 now you can solve it like any other equation first. pretend the bases aren't there. 2(x+1) = 3 2x+2 = 3 2x = 1 x = 0.5

OpenStudy (photon336):

\[9^{0.5+1} = 9^{\frac{ 3 }{ 2 }} = 27 \]

OpenStudy (photon336):

let's do it another way, have you learned about logs?

OpenStudy (vladthepro):

Yes, but still not that good at them yet

OpenStudy (photon336):

\[9^{3/2} = \sqrt[2]{9^{3}} = \sqrt{729} = 27\]

OpenStudy (photon336):

yeah so let's do it the other way \[9^{x+1} = 27 \] First we apply the product rule of logs \[\log_{b}(x^{y}) = y*\log_{b}(x)\] Then we just solve it like any other equation \[(x+1)Log_{10}*9 = Log(27)\] \[xlog*9+\log_{10}9 = Log_{10}(27)\] \[\frac{ (Log_{10}27-\log_{10}9) }{ Log_{10}~9 } = \frac{ 1 }{ 2 }\]

OpenStudy (vladthepro):

Oh i see now. Thx for taking ur time to help me :)

OpenStudy (photon336):

yeah np